OpenStudy (anonymous):

How do you solve the equation 7^(7n+7)=2,401 ?

OpenStudy (anonymous):

Take a log on both sides. You get (7n+7)*log(7) = log(2401) 7n + 7 = log (2401)/log(7) Now solve for n.

OpenStudy (anonymous):

easier way to do: 7^(7n+7)=2,401 7^(7n+7)=7^4 compare 7n+7 with 4 7n+7=4 7n=4-7=-3 n=-3/7

OpenStudy (anonymous):

Well, that might be easier for this one problem. But, not the best approach for a more generic one. Taking "log" will always lead to the right answer.

OpenStudy (anonymous):

yes :) i agree, there are problems which need to use log to solve. but in this question using log is harder :)